# RaSeRec: Retrieval-Augmented Sequential Recommendation

### v2

@@ Abstract
Although prevailing supervised and self-supervised learning (SSL)-augmented sequential recommendation (SeRec) models have achieved improved performance with powerful neural network architectures, we argue that they still suffer from two limitations: (1) Preference Drift, where models trained on past data can hardly accommodate evolving user preference; and (2) Implicit Memory, where head patterns dominate parametric learning, making it harder to recall long tails. In this work, we explore retrieval augmentation in SeRec, to address these limitations. To this end, we propose a Retrieval-Augmented Sequential Recommendation framework, named RaSeRec, the main idea of which is to maintain a dynamic memory bank to accommodate preference drifts and retrieve relevant memories to augment user modeling explicitly. It consists of two stages: (i) collaborative-based pre-training, which learns to recommend and retrieve; (ii) retrieval-augmented fine-tuning, which learns to leverage retrieved memories. Extensive experiments on three datasets fully demonstrate the superiority  and effectiveness  of  RaSeRec [^1] The implementation code will be available at https://github.com/HITsz-TMG/RaSeRec..
RaSeRec: Retrieval-Augmented Sequential Recommendation
Xinping Zhao1, Baotian Hu1🖂, Yan Zhong2, Shouzheng Huang1, Zihao Zheng1, Meng Wang3, Haofen Wang3🖂, Min Zhang1 1Harbin Institute of Technology (Shenzhen), 2Peking University, 3Tongji University zhaoxinping@stu.hit.edu.cn, mengwangtj@tongji.edu.cn, carter.whfcarter@gmail.com, {hubaotian, zhangmin2021}@hit.edu.cn
footnotetext: 🖂Corresponding authors.
1 Introduction
Sequential Recommendation (SeRec) aims to predict the next items that users would like to adopt, by calculating the matching probabilities between their historical interactions and the candidate items Kang and McAuley (2018); Zhao et al. (2024a); Hidasi et al. (2016). As such, it is essential to learn high-quality user representations from the user’s historical interactions for better recommendation performance Xie et al. (2022); Qiu et al. (2022); Liu et al. (2021). However, with enormous parameters to be optimized, existing SeRec models commonly suffer from the data sparsity issue, i.e., users only interacted with an extremely small fraction of the total items Isinkaye et al. (2015), which makes it difficult to learn high-quality user representations.
@@ 1 Introduction
Recently, there has been a surge of research interest in leveraging self-supervised learning (SSL) within the realm of SeRec to alleviate the data sparsity issue caused by sparse supervised signals Xie et al. (2022); Qiu et al. (2022); Liu et al. (2021); Hao et al. (2024); Dang et al. (2023). Generally, it involves two procedures: (1) data augmentation, which generates multiple views for each user sequence, and (2) self-supervised learning, which maximizes the agreement between different views of the same (or similar) user sequence. In this way, SSL-Augmented models endow their model backbones with additional parametric knowledge to generate more high-quality user representations than those without SSL Wu et al. (2021).
Figure 1: Comparison between Vanilla, SSL-Augmented, and Retrieval-Augmented SeRec paradigms, where their system working flows are illustrated from left to right, respectively. $s_{u}$ denotes the user sequence, while $s_{u}^{\prime},s_{u}^{\prime\prime}$ represent two augmented views. SeqRec denotes the  sequence  encoder, referring  to §2 for more  technical  details.
Despite their effectiveness, we argue that current SSL-Augmented ones still suffer from some limitations:
@@
  • Preference Drift. In real-world recommender systems, user preference frequently drifts as time goes on Chen et al. (2022a); Lo et al. (2018) due to many reasons, e.g., life changes Aghdam et al. (2015). As such, models trained on past user preferences can hardly handle preference drifts and may recommend undesirable items to users, which leads to sub-optimal performance.
@@
  • Implicit Memory. Existing methods encode user sequential patterns into implicit memory (model parameters), where the distribution of observed patterns usually obeys a power law Clauset et al. (2009); Milojevic (2010). As such, long-tailed patterns that lack supervised signals may be overwhelmed by head ones during training, making it hard to recall long-tailed ones during inference.
@@
In this work, we focus on exploring Retrieval-Augmented Generation (RAG) Gao et al. (2023) in the SeRec, to address the above limitations. Though being well-studied in Large Language Models (LLMs) Lewis et al. (2020); Izacard and Grave (2021); Izacard et al. (2023); Zhao et al. (2024c); Asai et al. (2024), RAG is rarely explored in SeRec. The main idea of RAG is to retrieve relevant documents from an external knowledge base to facilitate LLMs’ generation, especially for knowledge-intensive tasks Lewis et al. (2020). For example, RAG Lewis et al. (2020) uses the input sequence as the query to retrieve text documents from Wikipedia and then adopts them as additional context to generate the target sequence; Self-RAG Asai et al. (2024) learns to retrieve passages on demand adaptively, and interleaves generation and retrieval with reflection tokens by unifying them as the next token prediction. Compared to SSL (similar to pre-training), RAG allows us to accommodate preference drift rapidly and recall long tails via looking up an external memory  bank  directly.
@@
Here, we wish to bring the RAG’s superiority into SeRec to enhance user representation learning, which differs from NLP tasks since they model language semantics while recommender systems model collaborative semantics Zheng et al. (2024a). To address the above limitations of SSL-Augmented SeRec models, we design a retrieval-augmented mechanism to improve the quality of user representations. Specifically, it consists of two key components: (1) memory retrieval, which recalls collaborative memories for the input user, and (2) representation augmentation, which leverages retrieved memories to augment user representation. On the one hand, memory retrieval adapts well to dynamic preference drifts with a real-time update memory bank. On the other hand, representation augmentation explicitly leverages retrieved memories, rendering it easier  to  recall  long  tails.
Figure 1 illustrates the difference between the three SeRec paradigms as well as the continuation between them. The vanilla SeRec is trained upon supervised signals to predict the next item based on the input user sequence Fang et al. (2020). The SSL-Augmented SeRec further enriches the parametric knowledge of the model backbone with SSL signals Yu et al. (2024). Standing on their shoulders, retrieval-augmented SeRec handles dynamic preference drifts and relieves the heavy reliance on implicit memory by memory retrieval and representation augmentation. Note that the retrieval-augmented SeRec is designed to further perfect existence, rather than completely  replacing  them.
@@
To this end, we propose a novel Retrieval-Augmented (RA) Sequential Recommendation paradigm, named RaSeRec, which learns to refine user representations with retrieved memories. RaSeRec’s training process consists of two key stages: (i) collaborative-based pre-training and (ii) retrieval-augmented fine-tuning. In the collaborative-based pre-training stage, the model is trained with two objectives built upon collaborative signals: (1) recommendation learning, which learns to predict the next item based on input user sequence, and (2) retrieval training, which learns to retrieve memories with the same (or similar) preference as the input user sequence. During retrieval-augmented fine-tuning, we treat <user sequence, target item> pairs as the reference set except the current input user sequence. We then use the pre-trained model to encode the reference set into a memory bank. Given a user sequence, our model first retrieves similar collaborative memories from the memory bank. Then, a Retrieval-Augmented Module (RAM) is built upon the pre-trained encoder, which learns to augment the representation of the input user with retrieved memories. An intuitive explanation of our idea is that it can be seen as an open-book exam. Specifically, with retrieval augmentation, RaSeRec does not have to memorize all sequential patterns whereas it learns to use retrieved ones (cheat sheets) from the memory bank.
Note that RaSeRec is model-agnostic and can be applied to any ID-based SeRec backbones. Here, we adopt SASRec Kang and McAuley (2018) as the backbone since it is simple yet effective. We also implement RaSeRec on different backbones (see §4.3). In summary, our contributions are: (1) We unveil the major issues existing models suffer from and shed light on the potential of RAG in SeRec. (2) We propose a new learning paradigm, RaSeRec, which explicitly retrieves collaborative memories and then learns to use them for better user representation modeling. (3) Extensive experiments on three benchmark datasets fully demonstrate that RaSeRec improves overall performance as well as enhances long-tailed recommendations.
2 Preliminaries
2.1 Problem Formulation
@@
$v_{*}=\arg\underset{v_{i}\in\mathcal{V}}{\max}{P\left(v_{|s_{u}|+1}^{(u)}=v_{i}|s_{u}\right)}.$ (1)
where $v_{*}$ denotes the predicted next item with the largest matching probability. In what follows, we employ bold lowercase and uppercase symbols to represent vectors as well as matrices, respectively.
2.2 Transformer for SeRec
In this section, we introduce how to model users’ historical interactions to get their representations. Specifically, we adopt SASRec Kang and McAuley (2018) as the model backbone, whose encoding module is based on the Transformer encoder Vaswani et al. (2017). To leverage the transformer’s strong encoding ability, we first convert items to embeddings. Then, we apply the Transformer encoder to generate the user representation.
Embedding Layer. Formally, an embedding table ${\mathbf{V}}\in\mathbb{R}^{|\mathcal{V}|\times d}$ is created upon the whole item set $\mathcal{V}$ to project each item (one-hot representation) into a low-dimensional dense vector for better sequential modeling, where $d$ is the embedding dimension. Additionally, to be aware of the time order of the sequence, a learnable position encoding matrix $\mathbf{P}\in\mathbb{R}^{T\times d}$ is constructed, where $T$ is the maximum sequence length. Owing to the constraint of the maximum sequence length $T$, when inferring user $u$ representation at time step $t+1$, we truncate the user sequence $s_{u,t}$ to the last $T$ items if $t>T$:
$s_{u,t}=[v_{t-T+1},v_{t-T+2},...,v_{t}],$ (2)
@@ 2.2 Transformer for SeRec
${\mathbf{h}}_{i}^{0}={\mathbf{v}}_{i}+{\mathbf{p}}_{i},$ (3)
@@
Transformer Encoder. A Transformer encoder is composed of a multi-head self-attention module and a position-wise feed-forward Network. More technical details about the above two modules can be found in Kang and McAuley (2018) and Vaswani et al. (2017). We stack multiple blocks to capture more complex sequential patterns for high-quality user representation. Specifically, an ${L}$-layer Transformer encoder (unidirectional attention) is applied to refine the hidden representation of each item in $s_{u,t}$, which can be  formulated  as  follows:
${\mathbf{H}}^{L}=\rm{Trm}({\mathbf{H}}^{0}),$ (4)
@@
where $\rm{Trm}(\cdot)$ denotes the $L$-layer Transformer encoder, ${\mathbf{H}}^{L}=[\mathbf{h}_{t-T+1}^{L},...,\mathbf{h}_{t}^{L}]$ is the refined hidden representations of the sequence. The last refined hidden representation $\mathbf{h}_{t}^{L}$ is selected as the representative of this user sequence. Here, we omit the superscript $L$ and subscript $t$ for brevity, i.e., $\mathbf{h}$. In addition, we use $\mathrm{SeqEnc(\cdot)}$ to denote the sequence encoder, that  is  $\mathbf{h}=\mathrm{SeqEnc}(s_{u,t})$  for  simplicity.
@@
Figure 2: The overall system framework of the proposed RaSeRec. The upper layer illustrates the workflow of collaborative-based pre-training while the bottom layer shows the workflow of retrieval-augmented fine-tuning.
3 Methodology
@@
We propose the Retrieval-Augmented Sequential Recommendation (RaSeRec) paradigm, which endows SeRec to accommodate preference drift and recall long-tailed patterns, as shown in Figure 2. Specifically, it consists of a two-stage training strategy: (1) collaborative-based pre-training, which learns to recommend and retrieve based on collaborative signals (§3.1); (2) retrieval-augmented fine-tuning, which learns to leverage retrieved memories to augment user representation (§3.2). Lastly, we introduce the inference process and analyze the complexity of RaSeRec (§3.3). The notation table can be found in Appendix A. The learning algorithm of RaSeRec is provided  in  Appendix J.
3.1 Collaborative-based Pre-training
@@
In this section, we pre-train the model backbone $\mathrm{SeqRec}(\cdot)$ with two learning objections: (1) recommendation learning, which endows the backbone with the ability to generate the next item based on the input user sequence, and (2) retrieval training, which endows the backbone with the ability to retrieve memories with the same (or similar) preference as the input  user  sequence  representation.
@@
Recommendation Learning. After computing the user representation $\mathbf{h}\in\mathbb{R}^{d}$2.2), a prediction layer is built upon it to predict how likely user $u$ would adopt item $v_{i}$ at time step $t+1$. A simple yet effective solution is to calculate the inner product, i.e., $\mathbf{h}^{T}\textbf{v}_{i}$. After that, we employ the next item prediction over the whole item set $\mathcal{V}$ as the recommendation learning objective. To be specific, we compute the negative log-likelihood with the softmax as the recommendation learning objective:
$\mathcal{L}_{rec}=-\log\Big{(}e^{\mathbf{h}^{T}\mathbf{v}_{t+1}}/\sum_{v_{i}\in\mathcal{V}}e^{\mathbf{h}^{T}\mathbf{v}_{i}}\Big{)},$ (5)
@@
where $\mathbf{h},\mathbf{v}_{t+1},\mathbf{v}_{i}$ denote the last refined hidden representation $\mathbf{h}_{t}^{L}$, the embedding vector of the next item $v_{t+1}$, and the embedding vector of item $v_{i}\in\mathcal{V}$, respectively. We train the model to predict the next item auto-regressively similar to the language modeling’s objective Radford et al. (2018).

### v3

@@ Abstract.
Although prevailing supervised and self-supervised learning augmented sequential recommendation (SeRec) models have achieved improved performance with powerful neural network architectures, we argue that they still suffer from two limitations: (1) Preference Drift, where models trained on past data can hardly accommodate evolving user preference; and (2) Implicit Memory, where head patterns dominate parametric learning, making it harder to recall long tails. In this work, we explore retrieval augmentation in SeRec, to address these limitations. Specifically, we propose a Retrieval-Augmented Sequential Recommendation framework, named RaSeRec, the main idea of which is to maintain a dynamic memory bank to accommodate preference drifts and retrieve relevant memories to augment user modeling explicitly. It consists of two stages: (i) collaborative-based pre-training, which learns to recommend and retrieve; (ii) retrieval-augmented fine-tuning, which learns to leverage retrieved memories. Extensive experiments on three datasets fully demonstrate the superiority and effectiveness of RaSeRec. The implementation code is available at https://github.com/HITsz-TMG/RaSeRec.
Sequential Recommendation, Retrieval Augmentation, Preference Drift, Implicit Memory
Baotian Hu and Haofen Wang are the corresponding authors
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Recommender systemsccs: Information systems Information retrieval
1. Introduction
@@
Sequential Recommendation (SeRec) aims to predict the next items that users would like to adopt, by calculating the matching probabilities between their historical interactions and the candidate items (Kang and McAuley, 2018; Zhao et al., 2024a; Hidasi et al., 2016). As such, it is essential to learn high-quality user representations from the user’s historical interactions for better recommendation performance (Xie et al., 2022; Qiu et al., 2022; Liu et al., 2021). However, with enormous parameters to be optimized, existing SeRec models commonly suffer from the data sparsity issue, i.e., users only interacted with a tiny fraction of the total items (Isinkaye et al., 2015), making it difficult to learn high-quality user representations. Recently, there has been a surge of research interest in leveraging self-supervised learning (SSL) within the realm of SeRec to alleviate the data sparsity issue caused by sparse supervised signals (Xie et al., 2022; Qiu et al., 2022; Liu et al., 2021; Hao et al., 2024; Dang et al., 2023). Generally, it involves two procedures: (1) data augmentation, which generates multiple views for each user sequence, and (2) self-supervised learning, which maximizes the agreement between different views of the same (or similar) user sequence. In this way, SSL-Augmented models endow their model backbones with additional parametric knowledge to generate more high-quality user representations than those without SSL (Wu et al., 2021). Despite their effectiveness, we argue that current SSL-Augmented ones still suffer from some severe limitations:
@@
  • Preference Drift. In real-world recommender systems, user preference frequently drifts as time goes on (Chen et al., 2022a; Lo et al., 2018) due to many reasons, e.g., life changes (Aghdam et al., 2015). As such, models trained on past user preferences can hardly handle preference drifts and may recommend undesirable items to users, which leads to sub-optimal performance.
@@
  • Implicit Memory. Existing methods encode user sequential patterns into implicit memory (model parameters), where the distribution of observed patterns usually obeys a power law (Clauset et al., 2009; Milojevic, 2010). As such, long-tailed patterns that lack supervised signals may be overwhelmed by head ones during training, making it hard to recall long-tailed ones during inference.
@@
In this work, we explore Retrieval-Augmented Generation (RAG) (Gao et al., 2023) in SeRec, to address the above limitations. Though being well-studied in Large Language Models (LLMs) (Lewis et al., 2020; Izacard and Grave, 2021; Izacard et al., 2023; Zhao et al., 2024c; Asai et al., 2024), RAG is rarely explored in SeRec. The main idea of RAG is to retrieve relevant documents from an external knowledge base to facilitate LLMs’ generation, especially for knowledge-intensive tasks (Lewis et al., 2020). For example, RAG (Lewis et al., 2020) uses the input sequence as the query to retrieve text documents from Wikipedia and then adopts them as additional context to generate the target sequence; Self-RAG (Asai et al., 2024) learns to retrieve passages on demand adaptively, and interleaves generation and retrieval with reflection tokens by unifying them as the next token prediction. Compared to SSL (similar to pre-training), RAG allows us to accommodate preference drift rapidly and recall long tails via looking up an external memory bank directly.
@@
Here, we wish to bring the RAG’s superiority into SeRec to enhance user representation learning, which differs from NLP tasks since they model language semantics while recommender systems model collaborative semantics (Zheng et al., 2024b). To address the above limitations of SSL-Augmented SeRec models, we design a retrieval-augmented mechanism to improve the quality of user representations. Specifically, it consists of two key components: (1) memory retrieval, which recalls collaborative memories for the input user, and (2) representation augmentation, which leverages retrieved memories to augment user representation. On the one hand, memory retrieval adapts well to dynamic preference drifts with a real-time update memory bank. On the other hand, representation augmentation explicitly leverages retrieved memories, rendering it easier to recall long tails. Figure 1 illustrates the difference between the three SeRec paradigms and the continuation between them. The vanilla SeRec is trained on supervised signals to predict the next item based on the input user sequence (Fang et al., 2020). The SSL-Augmented SeRec further enriches the parametric knowledge of the model backbone with SSL signals (Yu et al., 2024). Standing on their shoulders, the retrieval-augmented SeRec handles dynamic preference drifts and relieves the heavy reliance on implicit memory by memory retrieval and representation augmentation. Note that the retrieval-augmented SeRec is designed to further perfect existence, rather than completely replacing them.
Figure 1. Comparison between Vanilla, SSL-Augmented, and Retrieval-Augmented SeRec paradigms, where their system working flows are illustrated from left to right, respectively. $s_{u}$ denotes the user sequence, while $s_{u}^{\prime},s_{u}^{\prime\prime}$ represent two augmented views. SeqEnc denotes the sequence encoder, referring to §2.2 for more technical details.
@@
To this end, we propose a novel Retrieval-Augmented Sequential Recommendation paradigm, named RaSeRec, which learns to refine user representations with retrieved memories. Specifically, RaSeRec’s training process consists of two key stages: (i) collaborative-based pre-training and (ii) retrieval-augmented fine-tuning. In the collaborative-based pre-training stage, the model is trained with two objectives built upon collaborative signals: (1) recommendation learning, which learns to predict the next item based on input user sequence, and (2) retrieval training, which learns to retrieve memories with the same (or similar) preference as the input user sequence. During the retrieval-augmented fine-tuning stage, we treat ¡user sequence, target item¿ pairs as the reference set except the current input user sequence. We then use the pre-trained model to encode the reference set into a memory bank. Given a user sequence, our model first retrieves similar collaborative memories from the memory bank. Then, a Retrieval-Augmented Module (RAM) is built upon the pre-trained encoder, which learns to augment the representation of the input user with retrieved memories. An intuitive explanation of our idea is that it can be seen as an open-book exam. Specifically, with retrieval augmentation, RaSeRec does not have to memorize all sequential patterns whereas it learns to use retrieved ones (cheat sheets) from the memory bank. Note that RaSeRec is model-agnostic and can be applied to any ID-based SeRec backbones. Here, we adopt SASRec (Kang and McAuley, 2018) as the backbone since it is simple yet effective. We also implement it on different backbones (see §4.3). In summary, our contributions are:
  • We unveil the major issues existing models suffer from and shed light on the potential of RAG in SeRec.
  • We propose a new SeRec paradigm, RaSeRec, which explicitly retrieves collaborative memories and then learns to use them for better user representation modeling.
  • Extensive experiments on three benchmark datasets fully demonstrate that RaSeRec improves overall performance, enhances long-tailed recommendation, and alleviates preference drift.
2. Preliminaries
2.1. Problem Formulation
@@
(1) $v_{*}=\arg\underset{v_{i}\in\mathcal{V}}{\max}{P\left(v_{|s_{u}|+1}^{(u)}=v_{i}|s_{u}\right)},$
where $v_{*}$ denotes the predicted item with the largest matching probability. In what follows, we employ bold lowercase and uppercase symbols to represent vectors and matrices, respectively.
2.2. Transformer for SeRec
In this section, we introduce how to model users’ historical interactions to get their representations. Specifically, we adopt SASRec (Kang and McAuley, 2018) as the model backbone, whose encoding module is based on the Transformer encoder (Vaswani et al., 2017). To leverage the transformer’s strong encoding ability, we first convert items to embeddings. Then, we apply the Transformer encoder to generate the user representation.
2.2.1. Embedding Layer
Formally, an embedding table ${\mathbf{V}}\in\mathbb{R}^{|\mathcal{V}|\times d}$ is created upon the whole item set $\mathcal{V}$ to project each item (one-hot representation) into a low-dimensional dense vector for better sequential modeling, where $d$ is the embedding dimension. Additionally, to perceive the time order of the sequence, a learnable position encoding matrix $\mathbf{P}\in\mathbb{R}^{T\times d}$ is constructed, where $T$ is the maximum sequence length. Owing to the constraint of the maximum sequence length $T$, when inferring user $u$ representation at time step $t+1$, we truncate the user sequence $s_{u,t}$ to the last $T$ items if $t>T$, otherwise left unchanged:
(2) $s_{u,t}=[v_{t-T+1},v_{t-T+2},...,v_{t}],$
@@ 2.2.1. Embedding Layer
(3) ${\mathbf{h}}_{i}^{0}={\mathbf{v}}_{i}+{\mathbf{p}}_{i},$
@@ 2.2.2. Transformer Encoder
A Transformer encoder is composed of a multi-head self-attention module and a position-wise feed-forward Network. More technical details about these two modules can be found in (Kang and McAuley, 2018) and (Vaswani et al., 2017). We stack multiple blocks to capture more complex sequential patterns for high-quality user representation. Specifically, an ${L}$-layer Transformer encoder (unidirectional attention) is applied to refine the hidden representation of each item in ${\mathbf{H}}^{0}$, which can be formulated as follows:
(4) ${\mathbf{H}}^{L}=\rm{Trm}({\mathbf{H}}^{0}),$
@@
where $\rm{Trm}(\cdot)$ denotes the $L$-layer Transformer encoder, ${\mathbf{H}}^{L}=[\mathbf{h}_{t-T+1}^{L},...,\mathbf{h}_{t}^{L}]$ is the refined hidden representations of the sequence. The last refined hidden representation $\mathbf{h}_{t}^{L}$ is selected as the representative of the user sequence. Here, we omit the superscript $L$ and subscript $t$ for brevity, i.e., $\mathbf{h}$. In addition, we use $\mathrm{SeqEnc(\cdot)}$ to denote the sequence encoder, that is $\mathbf{h}=\mathrm{SeqEnc}(s_{u,t})$ for simplicity.
@@
Figure 2. The overall system framework of the proposed RaSeRec. The upper layer illustrates the workflow of collaborative-based pre-training while the bottom layer shows the workflow of retrieval-augmented fine-tuning.
3. Methodology
@@
We propose the Retrieval-Augmented Sequential Recommendation (RaSeRec) paradigm, which endows SeRec to accommodate preference drift and recall long-tailed patterns, as shown in Figure 2. Specifically, it consists of a two-stage training strategy: (1) collaborative-based pre-training, which learns to recommend and retrieve based on collaborative signals (§3.1); (2) retrieval-augmented fine-tuning, which learns to leverage retrieved memories to augment user representation (§3.2). Lastly, we introduce the inference process and analyze the complexity of RaSeRec (§3.3). The notation table can be found in Appendix A.
3.1. Collaborative-based Pre-training
@@
In this section, we pre-train the model backbone $\mathrm{SeqEnc}(\cdot)$ with two learning objections: (1) recommendation learning, which endows the backbone with the ability to generate the next item based on the input user sequence, and (2) retrieval training, which endows the backbone with the ability to retrieve memories with the same (or similar) preference as the input user sequence.
3.1.1. Recommendation Learning
@@
After computing the user representation $\mathbf{h}\in\mathbb{R}^{d}$2.2), a prediction layer is built upon it to predict how likely user $u$ would adopt item $v_{i}$ at time step $t+1$. A simple yet effective solution is to calculate the inner product, i.e., $\mathbf{h}^{T}\textbf{v}_{i}$. After that, we employ the next item prediction over the whole item set $\mathcal{V}$ as the recommendation learning objective. To be specific, we compute the negative log-likelihood with the softmax as the recommendation learning objective:
(5) $\mathcal{L}_{rec}=-\log\Big{(}e^{\mathbf{h}^{T}\mathbf{v}_{t+1}}/\sum_{v_{i}\in\mathcal{V}}e^{\mathbf{h}^{T}\mathbf{v}_{i}}\Big{)},$
@@
where $\mathbf{h},\mathbf{v}_{t+1},\mathbf{v}_{i}$ denote the last refined hidden representation $\mathbf{h}_{t}^{L}$, the embedding vector of the next item $v_{t+1}$, and the embedding vector of item $v_{i}\in\mathcal{V}$, respectively. We train the model to predict the next item auto-regressively similar to the language modeling’s objective (Radford et al., 2018).
3.1.2. Retrieval Training